# Harnessing PU Learning for Enhanced Cloud-based DDoS Detection: A Comparative Analysis

### v1

@@ Abstract.
Cloud Computing and Security, Cybersecurity, Distributed Denial-of-Service (DDoS), Machine Learning, Positive-unlabeled (PU) Learning, Anomaly Detection, XGBoost, Random Forest, Support Vector Machine, Naïve Bayes
copyright: Noneccs: Computer systems organization Cloud computingccs: Networks Denial-of-service attacksccs: Computing methodologies Anomaly detection
@@ 2.2.3. Class Prior Incorporation
This technique modifies standard learning methods by applying mathematics from the Selected Completely At Random (SCAR) assumption, utilizing the provided class prior.
@@ 2.4. Assumptions and Limitations
  • The assumption that all unlabeled examples are negative may lead to a “close-world assumption,” which, while simplistic, can be serviceable in certain contexts.
@@ 3.1. Anomaly Detection in Cloud Infrastructure
By training a PU learning model on known “positive” (malicious or anomalous) examples and “unlabeled” (potentially benign or normal) data, the system can learn to flag potential security threats or system malfunctions effectively.
@@ 3.2. Vulnerability Detection
PU learning can aid in proactively identifying vulnerabilities in cloud-based applications, services, or infrastructure. The model can be trained on known “positive” examples of vulnerabilities and “unlabeled” data, enabling security teams to address potential weaknesses before they can be exploited.
@@ 17 DDoS attack scenarios
    Managing Extreme Values: As part of the preprocessing, we also need to address any extreme values in the dataset that may skew results or lead to errors during model training. We define a threshold for what constitutes an “extreme” value. For any numeric entry that exceeds this threshold, we replace the value with a hashed representation of the original value. This transformation helps maintain numerical stability while ensuring that extreme values do not negatively impact the learning algorithms.
@@ 5.2. PU Learning Implementation
find_first_csv(directory) Searches the specified directory for the first CSV file, returning the path or None if no file is found.
main() Main function that orchestrates the loading of the dataset, preprocessing, model training, and evaluation.

### v2

@@ Abstract.
Cloud Computing and Security, Cybersecurity, Distributed Denial-of-Service (DDoS), Machine Learning, Positive-unlabeled (PU) Learning, Negative-Unlabeled (NU) Learning, Anomaly Detection, XGBoost, Random Forest, Support Vector Machine, Naïve Bayes
copyright: None
@@ 1.2. Paper Structure
  • Section 8 reframes the task of DDoS detection as an NU Learning problem and analyzes the performance of the original PU configuration against the reconceptualized NU setup.
@@ 2.2.3. Class Prior Incorporation
This technique modifies standard learning methods by applying mathematics from the Selected Completely At Random (SCAR) assumption–a sampling method where individuals or items are selected randomly from a population without any specific criteria or bias–utilizing the provided class prior.
@@ 2.4. Assumptions and Limitations
[…]
@@ 3.1. Anomaly Detection in Cloud Infrastructure
[…]
@@ 3.2. Vulnerability Detection
[…]
@@ 17 DDoS attack scenarios
[…]
@@ 5.2. PU Learning Implementation
evaluate_pu_model(model_class, x_train, y_train, x_test, y_test, hold_out_ratio) Evaluates a Positive-Unlabeled (PU) learning model by fitting it to the training data, making predictions, and printing evaluation metrics.
main() Main function that orchestrates the loading of the dataset, preprocessing, model training, and evaluation. Allows toggling of evaluations for base ML models and PU learning models.
@@ 7.6. Approach Determination
8. Contrastive Study of PU Learning and NU Learning Approaches in DDoS Detection
In this section, we conduct a comparative analysis between the PU Learning and NU Learning approaches for Cloud-based DDoS detection. By analyzing the differences in how each framework handles the labeling of data and the impact of these strategies on model performance, we aim to highlight the strengths and limitations of each approach.
8.1. Preprocessing and Binary Label Assignment
In our preprocessing stage, we assign a binary label to each data instance based on its prior classification as Benign or !Benign. If the prior data is labeled as benign, it receives a value of 0; otherwise, it receives a value of 1. This binary labeling system enables the classification of data into two distinct categories, which is foundational to the PU Learning framework that we apply in our problem setup.
8.2. PU Learning Problem Setup
Our problem setup is structured within the framework of PU Learning (Positive-Unlabeled Learning). In this context, instances labeled as !Benign represent the positive class (value of 1), while the benign instances are treated as the unlabeled class (value of 0). PU Learning is particularly suitable for scenarios in which only a subset of data instances is confidently labeled as positive, while the rest remain unlabelled rather than explicitly negative.
Our PU Learning model implementation follows the two-step approach detailed earlier in this paper. First, the model trains on the positive samples (labeled as !Benign) and treats the unlabeled samples as potential positive examples. In the second step, it refines the model to better distinguish between true positives and false positives among the unlabeled data. This approach is crucial in instances where data is asymmetrically labeled, as is the case with our dataset of interest: BCCC-cPacket-Cloud-DDoS-2024.
8.3. NU Learning: Concept and Setup
[…]
The implementation of NU Learning involves a similar two-step approach, wherein the model trains on the available negative (benign) samples and progressively learns to distinguish between benign and potentially non-benign instances. NU Learning is applied to understand how well the model can classify an instance based on its alignment with known benign features.
8.4. Extending the Comparative Analysis
To deepen our analysis, we extend the study to compare the efficacy of framing Cloud-based DDoS detection as either a PU Learning or NU Learning problem. By examining both setups, we can assess which approach more accurately captures the nuanced patterns in the dataset and thus enables more effective detection of anomalous (or potentially malicious) instances.
8.5. Dataset Modification for NU Learning
To implement NU Learning, we modify the dataset by flipping the binary label values in the last column. Every instance of a 0 in the last column is changed to a 1 and vice versa. After this modification, we output the count of rows ending in 0 and 1 both before and after the transformation to confirm the accuracy of the modification. We then re-run our analysis code with the modified dataset to generate a new bar chart for the NU Learning approach, allowing for a direct comparison between the PU and NU Learning outcomes.
Figure 7. NU Learning Problem Recontextualization and Post Processing
8.6. NU Learning Results
Figure 8. NU Learning Trial 1
Figure 9. NU Learning Trial 2
Figure 10. Comparison of Machine Learning (ML) Model Performance Metrics for DDoS Detection using NU Learning
Model $F_{1}$-Score ROC AUC Recall Precision
XGBoost 0.986158 0.987433 0.980350 0.992036
Random Forest 0.981666 0.983702 0.976302 0.987089
SVM 0.651965 0.676619 0.779440 0.560325
Naïve Bayes 0.633777 0.608283 0.956374 0.473918
Table 3. Results of the NU-Learning Approach
Figure 11. Evaluating Machine Learning Model Performance Metrics for DDoS Detection: PU Learning vs. NU Learning Problem Framing
In terms of the $F_{1}$-Score (as exhibited in Figure 11), both PU and NU Learning show consistent performance, with the exception of NU Learning’s Naïve Bayes implementation, which achieves a slightly higher score. When considering Recall, NU Learning surpasses PU Learning, particularly with its Naïve Bayes and Support Vector Machine (SVM) models. The performance of Random Forest and XGBoost, however, remains nearly identical across both learning frameworks. Similarly, the ROC AUC scores for both models are virtually the same. On the other hand, PU Learning excels in Precision when paired with Naïve Bayes and SVM, while Random Forest and XGBoost once again demonstrate comparable results in both setups.
8.7. Assessment Deliberation
The comparison of PU and NU Learning results shows that both approaches yield high performance with the XGBoost and Random Forest algorithms. However, XGBoost consistently achieves higher precision and recall scores across both setups, indicating it is particularly effective for identifying the nuanced distinctions in DDoS attack patterns. Meanwhile, SVM and Naïve Bayes underperform in both settings, suggesting that these algorithms may be less suitable for this problem due to their lower adaptability to the complex feature sets associated with DDoS detection.
As an aside, it is infeasible for us to incorporate prior knowledge into the Naïve Bayes implementation, which may account for its poorer performance. Similarly, the use of a linear support vector machine (SVM) likely contributes to SVM’s limited efficacy in this context.
Based on these findings, the PU Learning approach with the XGBoost algorithm is recommended for implementation, as it provides a more balanced and reliable model for distinguishing between benign and malicious traffic in Cloud-based DDoS detection tasks.